OpenStudy (anonymous):

5g + 4 (-5 + 3g) = 1 - g SHOW WORK PLEASE

OpenStudy (anonymous):

5g - 20 + 12g = 1 - g 6g - 20 + 12g = 1 18g -20 = 1 18g = 21

OpenStudy (anonymous):

THANK YOU! :)